White Mountain Packaging is evaluating the laser tag center project. During year 1, the laser tag center project is expected to have relevant revenue of 949,800 dollars, relevant variable costs of 206,500 dollars, and relevant depreciation of 77,600 dollars. In addition, White Mountain Packaging would have one source of fixed costs associated with the laser tag center project. White Mountain Packaging just signed a deal with Blue Eagle Media to develop an advertising campaign for use in the project. The terms of the deal require White Mountain Packaging to pay Blue Eagle Media either 123,800 dollars in 1 year if the project is pursued or 149,000 dollars in 1 year if the project is not pursued. Relevant net income for the laser tag center project in year 1 is expected to be 312,840 dollars. What is the tax rate expected to be in year 1?
Answer #1

The terms of the deal require White Mountain Packaging to pay Blue Eagle Media either 123,800 dollars in 1 year if the project is pursued or 149,000 dollars in 1 year if the project is not pursued.

Since there is expected to be a net income in year 1, one can safely make out that the project will be pursued. Hence, we consider $ 123,800 as the fixed cost for year 1.

Let T be the tax rate for year 1.

Earnings before taxes = Net income / (1 - T) = 312,840 / (1 - T)

Earnings before taxes = Revenue - variable cost - fixed cost - depreciation - interest = 949,800 - 206,500 - 123,800 - 77,600 - 0 =  541,900

Hence,  541,900 = 312,840 / (1 - T)

Hence, tax rate = T = 1 - 312,840 /  541,900 = 42.27%